the transition in hydrogen atom which will give rise to least energetic photon is                                                                    1)n=5 to n=4
2)n=4 to n=2
3)n=3 to n=1
4)n=2 to n=1                                                                              give me reason.

Grade:11

1 Answers

Khimraj
5 years ago
According to Bohr's theory , 
\Delta E\propto[1/n₁² - 1/n₂²] 
(a) n = 5 to n = 4 
[1/4² - 1/5²] = 1/16 - 1/25 = 9/400
(b) n = 4 to n = 2 
[1/2² - 1/4²] = 1/4 - 1/16 = 3/16
(c) n = 3 to n = 1
[1/1² - 1/3²] = 1 - 1/9 = 8/9 
(d) n = 2 to n = 1 
[1/1² - 1/2²] = 1 - 1/4 = 3/4
Here least value of [1/n₁² - 1/n₂²] for option (a)
Hope it clears. If you like answer then please approve it.